40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jan 20 20:20:36 2020, gautam, Update, PSL, PMC servo checkout PMCsweep.pdfOLTFmeas.pdf
    Reply  Wed Jan 22 11:22:39 2020, gautam, Update, PSL, PMC modulation depth measurement modDepth.pdf
       Reply  Wed Jan 22 20:12:36 2020, gautam, Update, PSL, PMC demodulator electrical characterization demodChar.pdfmixerChar.pdf
          Reply  Thu Jan 23 14:37:05 2020, gautam, Update, PSL, PMC VGA chip damaged? VGAchar.pdf
             Reply  Thu Jan 23 16:37:14 2020, rana, Update, PSL, PMC VGA chip damaged? 
             Reply  Thu Jan 23 22:10:01 2020, gautam, Update, PSL, PMC servo pulled out 
                Reply  Fri Jan 24 15:42:08 2020, gautam, Update, PSL, PMC servo restored 
                   Reply  Sun Jan 26 13:47:00 2020, gautam, Update, PSL, PMC servo characterization elecTFs.pdfVGAchar_postFix.pdfVGAlinearity_postFix.pdfnewOLTFs.pdf
                      Reply  Tue Jan 28 14:33:24 2020, gautam, Update, PSL, Inferred free-running frequency noise inLoopNoise_IMClocked.pdffreqNoiseComparison.pdf
          Reply  Thu Jan 23 18:52:31 2020, gautam, Update, PSL, PMC RFPD characterization PDresp.pdf
Message ID: 15139     Entry time: Wed Jan 22 11:22:39 2020     In reply to: 15135     Reply to this: 15143
Author: gautam 
Type: Update 
Category: PSL 
Subject: PMC modulation depth measurement 

Summary:

I estimate the PMC servo modulation depth to be approximately 50 mrad. This is only 15% lower than what was measured in Jan 2018, and cannot explain the ~x50 reduction of optical gain measured earlier in this thread. Later in the day, I also confirmed that the LO input to the ZAD-6 mixer is +7 dBm. So the crystal is not to blame.

Details:

  1. PSL frequency is locked to the IMC length.
  2. Arm lengths are locked to the PSL frequency using POX/POY.
  3. EX green laser locked to the X arm length using end PDH servo. GTRX was ~0.4 in this measurement, which is the nominal value.
  4. The 20dB coupled port of the beat between the EX and PSL lasers was monitored using the AG4395A in "Spectrum" units.
  5. The beat was set at ~90 MHz, and a spectrum was taken for ~100 MHz span centered at the beat frequency.
  6. The modulation depth is estimated by considering the ratio of power at the beat frequency relative to that 35.5 MHz away. See Attachment #1.

Assuming a finesse of 700 for the PMC, we expect an optical gain of 2*Pin*J0(50e-3)*J1(50e-3)/fp  ~ 1.2e-7 W/Hz (=0.089 GW/m). I can't find a measurement of the PMC RFPD transimpedance to map this onto a V/Hz value. 

Attachment 1: modDepth.pdf  127 kB  | Hide | Hide all
modDepth.pdf
ELOG V3.1.3-